express as the cofunction of a positive acute angle not using a calculator of sec -160 degrees.
Answer by venugopalramana(3286)

SEC(-160)=SEC(160)...SINCE SEC IS POSITIVE IN IV QUADRANT..SEC(-X)=SEC(X)
=SEC(180-20)=-SEC(20)...SINCE SEC IS NEGATIVE IN II QUADRANT..SEC(180-X)=-SEC(X)